OpenStudy (anonymous):

Find the inverse function for: f(x)=(√(1-x^2))/(x+1)

OpenStudy (anonymous):

I know the answer but I don't know how to get it.

OpenStudy (anonymous):

change f(x) to x and x to y. then isolate y in one side.

OpenStudy (anonymous):

Yeah I know that much but then I get stuck on \[(x(y+1))^2=1-y^2 \]

OpenStudy (xapproachesinfinity):

\(\Large \rm\color{teal}{is~that~function~like~this\\ f(x)=\frac{\sqrt{1-x^2}}{x+1}}\)

OpenStudy (xapproachesinfinity):

??

OpenStudy (xapproachesinfinity):

I'm gonna assume that's the function and continue First thing you need to do is to check for one-to-one property if it is satisfied in this case it is just looking at the function there is no need for restriction but better take a look at the graph https://www.desmos.com/calculator surely the horizontal line test passes

OpenStudy (xapproachesinfinity):

https://www.desmos.com/calculator/kgr0o4rk2k

OpenStudy (xapproachesinfinity):

Okay now the main part \(\large\bf {f(x)=\frac{\sqrt{1-x^2}}{x+1}}\) set is like this \(\large\bf {y=\frac{\sqrt{1-x^2}}{x+1}}\) change x and y places \(\large\bf {x=\frac{\sqrt{1-y^2}}{y+1}}\) Solve for y \(\large\bf {x(y+1)=\sqrt{1-y^2}}\) square both sides \(\large\bf {x^2(y+1)^2=1-y^2}\) \(\large\bf {x^2(y+1)^2-\color{red}{(1-y^2)}=(1-y^2)-\color{red}{(1-y^2)}}\) \(\large\bf {x^2(y+1)^2-\color{red}{(1-y^2)}=0}\) \(\large\bf {x^2(y+1)^2-(1-y)(1+y)=0}\) \(\large\bf {[x^2(y+1)-(1-y)](1+y)=0}\) \(\large\bf {[x^2y+x^2-1+y)](1+y)=0}\) we don't care about 1+y=0 our focus on the one that has x \(\large\bf {[(x^2+1)y+x^2-1)]=0}\) \(\large\bf {(x^2+1)y=1-x^2}\) \(\large\bf {y=\frac{1-x^2}{1+x^2}}\) then \(\large\bf {f^{-1}(x)=\frac{1-x^2}{1+x^2}}\)

OpenStudy (xapproachesinfinity):

Now check let's see if that's true. we don't want errors

OpenStudy (xapproachesinfinity):

hey you there?

OpenStudy (anonymous):

yeah sorry.. that's the right answer. Thank you sooo much!!

OpenStudy (xapproachesinfinity):

Okay! how do you know it is right? we have to check \(\large\bf {f(f^{-1}(x))=x}\) the best think to do is choose a value and evaluate we don't want to spend alot of time simplifying the expression \(\large\bf {f(f^{-1}(2))=2}\) we need to verify this

OpenStudy (xapproachesinfinity):

i checked and it is true^_^

OpenStudy (anonymous):

I'll get back to you in about an hour.. i got to go somewhere.. thank you so much anyways.

OpenStudy (xapproachesinfinity):

Alright! you are welcome make sure to check the steps^_^